OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 3/8 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are dimensions of the brand name television? Answer 32 inches by 64 inches 28 inches by 56 inches 7.5 inches by 15 inches 4.5 inches by 9 inches

OpenStudy (phi):

if g= 3/8 b then b= 8/3 g

OpenStudy (anonymous):

= ( 8/3)* (12 inches by 24 inches ) = 32 in by 64 in

OpenStudy (anonymous):

thank you : )

OpenStudy (anonymous):

Do you have any question?